# An exact dimension-reduced dynamic theory for developable surfaces and curve-fold origami

### v1

@@ Abstract
Curve-fold origami, composed of developable panels joined along a curved crease, exhibits rich dynamic behaviors relevant to metamaterials and soft robotic systems. Despite multiple approximated models (shell model, Kirchhoff rod model, etc.), a comprehensive and exact dynamical theory for curve-fold origami remains absent, limiting the precise predictions of its dynamics, especially for those with wide panels. In this work, we develop an exact dimension-reduced theory that focuses on the dynamics of curve-fold origami, utilizing the intrinsic one-dimensional nature of developable surfaces. As is well known in differential geometry, a curve on a developable surface determines the local frame and the shape of the developable surface under isometric constraint, making the problem indeed one-dimensional. Starting from a single developable surface, we investigate the kinematics and kinetic energy of a moving developable surface. By overcoming the difficulty of describing the motion of local frames, we derive the exact velocity field of wide surfaces solely described by the motion of the reference curve, which leads to the kinetic energy of the entire surface. The elastic energy follows the Wunderlich functional applicable to wide panels. Owing to the one-dimensional feature, the Lagrangian of the system, composed of both kinetic and elastic energy, is a functional of the reference curve. Thus, we may variate the Lagrangian and derive a nonlinear dynamical theory for the reference curve, which comprises governing equations similar to the rod model but can precisely describe the motion of developable surfaces. The theory is validated consistently in both Lagrangian and Eulerian frameworks and is further extended to curved-fold origami modeled as a coupled bi-rod system. Utilizing our exact 1D model, we theoretically analyze the dynamical behaviors of various developable structures, revealing that the coupling of curvature and torsion along with the motion of local frames in our theory leads to the accurate modeling of arbitrarily deformed developable surfaces, which are validated by finite element analysis quantitatively. Our exact dimensional-reduced dynamical theory captures the motions of developable surfaces and curve-fold origami (with wide panels) accurately at relatively low computational cost (since it is 1D), offering a potential contribution to the design of dynamical developable structures.
@@ 1 Introduction
Origami, a traditional art form, enables complex three-dimensional shape transformations through prescribed crease patterns, and also inspires diverse engineering applications. Over the past centuries, efforts have been made by origamists, mathematicians and physicists on the geometry and statics of straight-fold origami (Izmestiev, 2017; Tachi, 2009; Demaine and Tachi, 2017; Liu and Paulino, 2017; Filipov et al., 2017), exploring the quasi-static transformations of origami under geometric constraints. More recently, these efforts have been expanded to the kinematics and dynamics of straight-fold systems, producing models that couple the geometric compatibility with elasticity to describe the motion of dynamic origami structures (Zou et al., 2024; Herkal et al., 2024). Such theories have facilitated the development of origami-based metamaterials (Silverberg et al., 2014; Liu et al., 2018) and robotic mechanisms (Rus and Sung, 2018). Beyond traditional straight folds, recently a novel branch of origami with curved folds, i.e., curve-fold origami, has gained more interest (Koschitz et al., 2008; Demaine et al., 2011). In addition to aesthetic design, curved folds offer unique mechanical capabilities—such as tunable stiffness, energy storage, and large shape change—that have motivated applications in soft robotics (Rus and Sung, 2018; Feng et al., 2024), metamaterials (Lee et al., 2021; Karami et al., 2024; Sun et al., 2024), architecture (Tachi and Epps, 2011; Mouthuy et al., 2012), and virtual-reality interfaces (Zhang et al., 2023). Besides, its extraordinary dynamic behaviors have enabled the designs of fast-develop robotic wings (Baek et al., 2020) and acoustic metamaterials (Zhai et al., 2020). Despite these advances, we notice that most of the dynamic theories rely on narrow ribbon or rod approximations, and an exact dynamical theory for curve-fold origami, in particular for those with wide panels, is still lacking, limiting further applications that require accurate predictions of dynamics at moderate computational cost. Addressing this gap is the primary objective of the present study.
@@
In contrast to the well-established geometric and static frameworks for developable surfaces and curved-fold origami, their dynamic behaviors and, in particular, the associated kinetic energy, remain poorly underdeveloped. Existing studies on dynamics are largely based on the dynamic Kirchhoff rod model for narrow and straight developable ribbons, according to previous works by Goriely et al. (2001), Dai et al. (2022) and Huang et al. (2024). Solving the Kirchhoff equations yields the motion of such structures. These models perform reasonably well for narrow ribbons (Yu et al., 2023; Shi et al., 2025). However, the Kirchhoff rod formulation does not enforce the isometric constraint of developable surfaces, leading to significant errors for highly twisted ribbons (Audoly and Neukirch, 2021), such as Möbius bands (Mahadevan and Keller, 1993; Starostin and van der Heijden, 2007). From the perspective of energy, this limitation arises because the line energy density of Kirchhoff rods scales as $\varepsilon_{crease}\in(A\kappa^{2}+B\tau^{2})$, which becomes inaccurate for configurations with large torsion (i.e. extremely curved case). In contrast, the Sadowsky functional exhibits the scaling $\varepsilon_{crease}\in(\kappa^{2}+\tau^{2})^{2}/\kappa^{2}$. To address this issue, Dias and Audoly (2014, 2015) develop a nonlinear rod model for narrow developable surfaces, employing constitutive laws derived from Sadowsky’s energy and enforcing the developability/isometric constraint. Despite these advances, current formulations remain limited to narrow ribbons and static settings. They do not provide a general dynamic theory applicable to developable surfaces of arbitrary width, which is required for large-scale applications such as space antennas.
@@ 2.2 Kinematics of developable surfaces
The first three terms in Eq. (19) represent the translational acceleration, rotational acceleration, centripetal acceleration, similar to the acceleration in a rigid body, while the fourth term is an acceleration induced by frame variations.
@@ 2.4.1 Constitutive laws of the effective rod
where $m_{i}$ is the component form of the internal moment $\mathbf{m}$. Notice that the geodesic curvature remains constant in isometric deformation, resulting in $\delta\omega_{3}=0$ and $m_{3}$ is thus a constraint moment irrelevant to the rods’ geometry. To ensure that the variations of the elastic energy for both systems are equal, the internal moment of the equivalent rod follows the constitutive law
$\mathbf{m}=\frac{1}{\kappa_{n}}\left(\frac{{\textrm{d}}}{{\textrm{d}}S}\frac{\partial\varepsilon}{\partial\eta^{\prime}}-\frac{\partial\varepsilon}{\partial\eta}\right)\mathbf{e}_{1}+\left(-\frac{\partial\varepsilon}{\partial\kappa_{n}}-\frac{\eta}{\kappa_{n}}(\frac{{\textrm{d}}}{{\textrm{d}}S}\frac{\partial\varepsilon}{\partial\eta^{\prime}}-\frac{\partial\varepsilon}{\partial\eta})\right)\mathbf{e}_{2}+m_{3}\mathbf{e}_{3},$ (27)
@@
to eliminate the boundary values in Eq. (25). In the narrow ribbon limit, i.e., $v_{0}\kappa_{g}\rightarrow 0$, the Wunderlich functional degenerates to Sadowsky (1930) functional, and Eqs. (27) are simplified into
@@
which are the same as Eq. (12) in (Dias and Audoly, 2014), indicating that our general rod model is consistent with the simplified one.
@@ 2.4.2 Dynamical equations
Substituting Eqs. (31) and (19) into (30), the full expression of the dynamical equation is
@@ 3.1 Lagrangian of curve-fold origami
$V_{bend+}+V_{bend-}=\Sigma_{\pm}\frac{1}{2}D\int\frac{\left(1+\eta_{\pm}^{2}\right)^{2}\kappa_{n\pm}^{2}}{\eta_{\pm}^{\prime}-\left(1+\eta_{\pm}^{2}\right)\kappa_{g}}\ln[1+\eta_{\pm}^{\prime}v_{0}-(1+\eta_{\pm}^{2})\kappa_{g}v_{0\pm}]{\textrm{d}}S,$ (40)
@@ 3.2 A nonlinear dynamic bi-rod model for curve-fold origami
which governs the cross-sectional orientation of the two equivalent rods. The Darboux frames are also shown in Fig. 4(a).
@@
$\mathbf{m}_{1}=\mathbf{m}_{inter}\cdot\mathbf{t}=\frac{\partial V_{fold}}{\partial\phi}=k_{c}(2\arctan(\frac{\kappa_{n}}{\kappa_{g}})-\phi_{0}).$ (45)
@@
$\mathbf{F}_{\pm}^{\prime}\pm\mathbf{F}_{inter}=\rho h[(v_{0\pm}+\frac{1}{2}v_{0\pm}^{2}\gamma_{\pm})\ddot{\mathbf{r}}_{0}+(\frac{1}{2}v_{0\pm}^{2}+\frac{1}{3}v_{0\pm}^{3}\gamma_{\pm})(\dot{\bm{\Omega}}_{\pm}\times\mathbf{l}_{\pm}+\bm{\Omega}_{\pm}\times(\bm{\Omega}_{\pm}\times\mathbf{l}_{\pm}))-\frac{1}{3}v_{0\pm}^{3}\dot{\eta}_{\pm}\bm{\Omega}_{\pm}^{\prime}\times\mathbf{l}_{\pm}],$ (46)
@@
$-\frac{1}{4}v_{0\pm}^{4}\dot{\eta}_{\pm}\mathbf{l}_{\pm}\times(\bm{\Omega}_{\pm}^{\prime}\times\mathbf{l}_{\pm})],$
$\mathbf{m}_{\pm}=\frac{1}{\kappa_{n\pm}}\left(\frac{{\textrm{d}}}{{\textrm{d}}S}\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}^{\prime}}-\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}}\right)\mathbf{e}_{1\pm}+\left(-\frac{\partial\varepsilon_{\pm}}{\partial\kappa_{n}}-\frac{\eta_{\pm}}{\kappa_{n\pm}}(\frac{{\textrm{d}}}{{\textrm{d}}S}\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}^{\prime}}-\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}})\right)\mathbf{e}_{2\pm}+m_{3\pm}\mathbf{e}_{3\pm},$
@@
$\mathbf{m}_{\Sigma}=\mathbf{m}_{e},$
@@
$\mathbf{F}^{\prime}_{\Sigma}=\Sigma_{\pm}\rho h[(\frac{1}{2}v_{0\pm}^{2}+\frac{1}{3}v_{0\pm}^{3}\gamma_{\pm})\mathbf{l}_{\pm}\times\ddot{\mathbf{r}}_{0}+(\frac{1}{3}v_{0\pm}^{3}+\frac{1}{4}v_{0\pm}^{4}\gamma_{\pm})\mathbf{l}_{\pm}\times(\dot{\bm{\Omega}}_{\pm}\times\mathbf{l}_{\pm}+\bm{\Omega}_{\pm}\times(\bm{\Omega}_{\pm}\times\mathbf{l}_{\pm})),$ (48)
@@
$-\frac{1}{4}v_{0\pm}^{4}\dot{\eta}_{\pm}\mathbf{l}_{\pm}\times(\bm{\Omega}_{\pm}^{\prime}\times\mathbf{l}_{\pm})],$
@@
$\mathbf{m}_{\pm}=\frac{1}{\kappa_{n\pm}}\left(\frac{{\textrm{d}}}{{\textrm{d}}S}\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}^{\prime}}-\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}}\right)\mathbf{e}_{1\pm}+\left(-\frac{\partial\varepsilon_{\pm}}{\partial\kappa_{n}}-\frac{\eta_{\pm}}{\kappa_{n\pm}}(\frac{{\textrm{d}}}{{\textrm{d}}S}\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}^{\prime}}-\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}})\right)\mathbf{e}_{2\pm}+m_{3\pm}\mathbf{e}_{3\pm}.$

### v2

@@ Abstract
Curve-fold origami, composed of developable panels joined along a curved crease, exhibits rich dynamic behaviors relevant to metamaterials and soft robotic systems. Despite multiple approximated models (shell model, Kirchhoff rod model, etc.), a comprehensive and exact dynamical theory for curve-fold origami remains absent, limiting the precise prediction of its dynamics, especially for those with wide panels. In this work, we develop an exact dimension-reduced theory that focuses on the dynamics of curve-fold origami, utilizing the intrinsic one-dimensional nature of developable surfaces. As is well known in differential geometry, a curve on a developable surface determines the local frame and the shape of the developable surface under isometric constraint, making the problem indeed one-dimensional. Starting from a single developable surface, we investigate the kinematics and kinetic energy of a moving developable surface. By overcoming the difficulty of describing the motion of local frames, we derive the exact velocity field of wide surfaces solely described by the motion of the reference curve, which leads to the kinetic energy of the entire surface. The elastic energy follows the Wunderlich functional applicable to wide panels. Owing to the one-dimensional feature, the Lagrangian of the system, composed of both kinetic and elastic energy, is a functional of the reference curve. Thus, we may variate the Lagrangian and derive a nonlinear dynamical theory for the reference curve, which comprises governing equations similar to the rod model but can precisely describe the motion of developable surfaces. The theory is validated consistently in both Lagrangian and Eulerian frameworks and is further extended to curved-fold origami modeled as a coupled bi-rod system. Utilizing our exact 1D model, we theoretically analyze the dynamical behaviors of various developable structures, revealing that the coupling of curvature and torsion along with the motion of local frames in our theory leads to the accurate modeling of arbitrarily deformed developable surfaces, which are validated by finite element analysis quantitatively. Our exact dimensional-reduced dynamical theory captures the motions of developable surfaces and curve-fold origami (with wide panels) accurately at relatively low computational cost (since it is 1D), offering a potential contribution to the design of dynamical developable structures.
@@ 1 Introduction
Origami, a traditional art form, enables complex three-dimensional shape transformations by folding along prescribed crease patterns, and also inspires diverse engineering applications. Over the past centuries, efforts have been made by origamists, mathematicians and physicists on the geometry and statics of straight-fold origami (Izmestiev, 2017; Tachi, 2009; Demaine and Tachi, 2017; Liu and Paulino, 2017; Filipov et al., 2017), exploring the quasi-static transformations of origami under geometric constraints. More recently, these efforts have been expanded to the kinematics and dynamics of straight-fold systems, producing models that couple the geometric compatibility with elasticity to describe the motion of dynamic origami structures (Zou et al., 2024; Herkal et al., 2024). Such theories have facilitated the development of origami-based metamaterials (Silverberg et al., 2014; Liu et al., 2018) and robotic mechanisms (Rus and Sung, 2018). Beyond traditional straight folds, recently a novel branch of origami with curved folds, i.e., curve-fold origami, has gained more interest (Koschitz et al., 2008; Demaine et al., 2011). In addition to aesthetic design, curved folds offer unique mechanical capabilities—such as tunable stiffness, energy storage, and large shape change—that have motivated applications in soft robotics (Rus and Sung, 2018; Feng et al., 2024), metamaterials (Lee et al., 2021; Karami et al., 2024; Sun et al., 2024), architecture (Tachi and Epps, 2011; Mouthuy et al., 2012), and virtual-reality interfaces (Zhang et al., 2023). Besides, its extraordinary dynamic behaviors have enabled the designs of fast-develop robotic wings (Baek et al., 2020) and acoustic metamaterials (Zhai et al., 2020). Despite these advances, we notice that most of the dynamic theories rely on narrow ribbon or rod approximations, and an exact dynamical theory for curve-fold origami, in particular for those with wide panels, is still lacking, limiting further applications that require accurate predictions of dynamics at moderate computational cost. Addressing this gap is the primary objective of the present study.
@@
In contrast to the well-established geometric and static frameworks for developable surfaces and curved-fold origami, their dynamic behaviors and, in particular, the associated kinetic energy, remain poorly underdeveloped. Existing studies on dynamics are largely based on the dynamic Kirchhoff rod model for narrow and straight developable ribbons, according to previous works by Goriely et al. (2001), Dai et al. (2022) and Huang et al. (2024). Solving the Kirchhoff equations yields the motion of such structures. These models perform reasonably well for narrow ribbons (Yu et al., 2023; Shi et al., 2025). However, the Kirchhoff rod formulation does not enforce the isometric constraint of developable surfaces, leading to significant errors for highly twisted ribbons (Audoly and Neukirch, 2021), such as Möbius bands (Mahadevan and Keller, 1993; Starostin and van der Heijden, 2007). From the perspective of energy, this limitation arises because the line energy density of Kirchhoff rods scales as $\varepsilon_{crease}\in(A\kappa^{2}+B\tau^{2})$, which becomes inaccurate for configurations with large torsion (i.e. extremely curved case). In contrast, the Sadowsky functional exhibits the scaling $\varepsilon_{crease}\in(\kappa^{2}+\tau^{2})^{2}/\kappa^{2}$. To address this issue, Dias and Audoly (2014, 2015) developed a nonlinear rod model for narrow developable surfaces, employing constitutive laws derived from Sadowsky’s energy and enforcing the developability/isometric constraint. Despite these advances, current formulations remain limited to narrow ribbons and static settings. They do not provide a general dynamic theory applicable to developable surfaces of arbitrary width, which is required for large-scale applications such as space antennas.
@@ 2.2 Kinematics of developable surfaces
The first three terms in Eq. (19) represent the translational acceleration, rotational acceleration, and centripetal acceleration, similar to the acceleration in a rigid body, while the fourth term is an acceleration induced by frame variations.
@@ 2.4.1 Constitutive laws of the effective rod
where $m_{i}$ is the component form of the internal moment $\mathbf{M}$, i.e., the moment of the part $(S,S_{0})$ acting on the other part $(0,S)$. Notice that the geodesic curvature remains constant in isometric deformation, resulting in $\delta\omega_{3}=0$ and $m_{3}$ is thus a constraint moment irrelevant to the rods’ geometry. To ensure that the variations of the elastic energy for both systems are equal, the internal moment of the equivalent rod follows the constitutive law
$\mathbf{M}=\frac{1}{\kappa_{n}}\left(\frac{{\textrm{d}}}{{\textrm{d}}S}\frac{\partial\varepsilon}{\partial\eta^{\prime}}-\frac{\partial\varepsilon}{\partial\eta}\right)\mathbf{e}_{1}+\left(-\frac{\partial\varepsilon}{\partial\kappa_{n}}-\frac{\eta}{\kappa_{n}}(\frac{{\textrm{d}}}{{\textrm{d}}S}\frac{\partial\varepsilon}{\partial\eta^{\prime}}-\frac{\partial\varepsilon}{\partial\eta})\right)\mathbf{e}_{2}+m_{3}\mathbf{e}_{3}.$ (27)
@@
to eliminate the boundary values in Eq. (25). In the narrow ribbon limit, i.e., $v_{0}\kappa_{g}\rightarrow 0$, the Wunderlich functional degenerates to the Sadowsky (1930) functional, and Eq. (27) is simplified into
@@
which is the same as Eq. (12) in (Dias and Audoly, 2014), indicating that our general rod model is consistent with the simplified one.
@@ 2.4.2 Dynamical equations
Substituting Eqs. (31) and (19) into (30), the full dynamical equation system is
@@ 3.1 Lagrangian of curve-fold origami
$V_{bend+}+V_{bend-}=\Sigma_{\pm}\frac{1}{2}D\int\frac{\left(1+\eta_{\pm}^{2}\right)^{2}\kappa_{n\pm}^{2}}{\pm\eta_{\pm}^{\prime}\mp\left(1+\eta_{\pm}^{2}\right)\kappa_{g}}\ln[1\pm\eta_{\pm}^{\prime}v_{0}\mp(1+\eta_{\pm}^{2})\kappa_{g}v_{0\pm}]{\textrm{d}}S,$ (40)
@@ 3.2 A nonlinear dynamic bi-rod model for curve-fold origami
which governs the cross-sectional orientation of the two equivalent rods. The Darboux frames are also shown in Fig. 4(a). According to Eq. (44), the angular velocity of the Darboux frame $\mathbf{\Omega}_{\pm}$ relates to the angular velocety of the Frenet frame $\mathbf{\Omega}_{f}$, following
$\mathbf{\Omega}_{\pm}=\mathbf{\Omega}_{f}\pm\frac{\dot{\phi}}{2}\mathbf{t}.$ (45)
@@
$\mathbf{M}_{inter}\cdot\mathbf{t}=\frac{\partial V_{fold}}{\partial\phi}=k_{c}(2\arctan(\frac{\kappa_{n}}{\kappa_{g}})-\phi_{0}).$ (46)
@@
$\mathbf{F}_{\pm}^{\prime}\pm\mathbf{F}_{inter}=\rho h[(v_{0\pm}+\frac{1}{2}v_{0\pm}^{2}\gamma_{\pm})\ddot{\mathbf{r}}_{0}+(\frac{1}{2}v_{0\pm}^{2}+\frac{1}{3}v_{0\pm}^{3}\gamma_{\pm})(\dot{\bm{\Omega}}_{\pm}\times\mathbf{l}_{\pm}+\bm{\Omega}_{\pm}\times(\bm{\Omega}_{\pm}\times\mathbf{l}_{\pm}))\mp\frac{1}{3}v_{0\pm}^{3}\dot{\eta}_{\pm}\bm{\Omega}_{\pm}^{\prime}\times\mathbf{l}_{\pm}],$ (47)
@@
$\mp\frac{1}{4}v_{0\pm}^{4}\dot{\eta}_{\pm}\mathbf{l}_{\pm}\times(\bm{\Omega}_{\pm}^{\prime}\times\mathbf{l}_{\pm})],$
$\mathbf{M}_{\pm}=\pm[\frac{1}{\kappa_{n\pm}}\left(\frac{{\textrm{d}}}{{\textrm{d}}S}\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}^{\prime}}-\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}}\right)\mathbf{e}_{1\pm}+\left(-\frac{\partial\varepsilon_{\pm}}{\partial\kappa_{n}}-\frac{\eta_{\pm}}{\kappa_{n\pm}}(\frac{{\textrm{d}}}{{\textrm{d}}S}\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}^{\prime}}-\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}})\right)\mathbf{e}_{2\pm}+m_{3\pm}\mathbf{e}_{3\pm}],$
@@
$\mathbf{M}_{\Sigma}=\mathbf{M}_{e}.$
@@
$\mathbf{F}^{\prime}_{\Sigma}=\Sigma_{\pm}\rho h[(\frac{1}{2}v_{0\pm}^{2}+\frac{1}{3}v_{0\pm}^{3}\gamma_{\pm})\mathbf{l}_{\pm}\times\ddot{\mathbf{r}}_{0}+(\frac{1}{3}v_{0\pm}^{3}+\frac{1}{4}v_{0\pm}^{4}\gamma_{\pm})\mathbf{l}_{\pm}\times(\dot{\bm{\Omega}}_{\pm}\times\mathbf{l}_{\pm}+\bm{\Omega}_{\pm}\times(\bm{\Omega}_{\pm}\times\mathbf{l}_{\pm}))\mp\frac{1}{3}v_{0\pm}^{3}\dot{\eta}_{\pm}\bm{\Omega}_{\pm}^{\prime}\times\mathbf{l}_{\pm}],$ (49)
@@
$\mp\frac{1}{4}v_{0\pm}^{4}\dot{\eta}_{\pm}\mathbf{l}_{\pm}\times(\bm{\Omega}_{\pm}^{\prime}\times\mathbf{l}_{\pm})],$
@@
$\mathbf{M}_{\pm}=\pm[\frac{1}{\kappa_{n\pm}}\left(\frac{{\textrm{d}}}{{\textrm{d}}S}\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}^{\prime}}-\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}}\right)\mathbf{e}_{1\pm}+\left(-\frac{\partial\varepsilon_{\pm}}{\partial\kappa_{n\pm}}-\frac{\eta_{\pm}}{\kappa_{n\pm}}(\frac{{\textrm{d}}}{{\textrm{d}}S}\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}^{\prime}}-\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}})\right)\mathbf{e}_{2\pm}+m_{3\pm}\mathbf{e}_{3\pm}].$